# Optimal observables for (non-)equilibrium quantum metrology from the master equation

### v1

@@ I Introduction
This paper is structured as follows. In section II we introduce our formalism to determine the SLD from the master equation. Section III applies the formalism to thermometry in the equilibrium steady state of quantum Brownian motion with a harmonic potential, before constructing in section IV the SLD for both temperature and relaxation rate in a non-equilibrium scenario based on a squeezed Gaussian state.
II Formalism
@@
III Equilibrium probe thermometry in Quantum Brownian Motion
@@
$\partial_{t}\hat{\rho}(t,T)=-i[\hat{H},\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\hat{\rho}\}]-2\gamma mT[\hat{x},[\hat{x},\hat{\rho}]],$ (11)
@@
$-\tfrac{1}{2}\left\langle\left\{\hat{A}_{i}\left[\hat{H},\hat{A}_{j}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{A}_{i},\left\{\hat{p},\left[\hat{A}_{j},\hat{x}\right]\right\}\right\}\right\rangle$
$-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{A}_{i},\left[\hat{x},\left[\hat{x},\hat{A}_{j}\right]\right]\right\}\right\rangle+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{A}_{i},\hat{A}_{j}\right\},\hat{x}\right]\right\}\right\rangle$
@@
$-i\gamma\left\langle\hat{A}_{j}\right\rangle\left\langle\left\{\hat{p},\left[\hat{A}_{i},\hat{x}\right]\right\}\right\rangle-\tfrac{\gamma T}{b}\left\langle\hat{A}_{j}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{A}_{i}\right]\right]\right\rangle$
@@
A careful evaluation of all $5\times 5$ entries of $M$ yields
$M=$ (21)
$\left[\begin{array}[]{ccccc}0&-2b\left\langle p^{2}\right\rangle&0&0&0\\[10.0pt] 2c\left\langle x^{2}\right\rangle&\begin{array}[]{c}2\gamma\left\langle p^{2}\right\rangle\\ -\frac{2\gamma T}{b}\end{array}&0&0&0\\[4.0pt] 0&0&0&2\gamma&-8b\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle-2b\\[4.0pt] 0&0&0&\begin{array}[]{c}8\gamma\left\langle p^{2}\right\rangle^{2}\\ -\frac{8\gamma T\left\langle p^{2}\right\rangle}{b}\end{array}&8c\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+2c\\[4.0pt] 0&0&\begin{array}[]{c}2b\\ +8c\left\langle x^{2}\right\rangle^{2}\end{array}&\begin{array}[]{c}-8b\left\langle p^{2}\right\rangle^{2}\\ -2c\end{array}&\begin{array}[]{c}-\frac{8\gamma T\left\langle x^{2}\right\rangle}{b}-2\gamma\\ +8\gamma\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle\end{array}\\ \end{array}\right]=\begin{bmatrix}0&-\frac{ba^{2}}{2}&0&0&0\\ \frac{a^{2}b}{2}&-\frac{\gamma a^{2}}{2}&0&0&0\\ 0&0&0&2\gamma&-\dfrac{a^{4}b^{2}}{2c}-2b\\ 0&0&0&-\dfrac{\gamma a^{4}}{2}&\frac{a^{4}b}{2}+2c\\ 0&0&\frac{a^{4}b^{2}}{2c}+2b&-\frac{a^{4}b}{2}-2c&-\frac{\gamma a^{4}b}{2c}-2\gamma\end{bmatrix},$ (37)
@@
Interesting structure emerges in the system of linear equations. Since the Caldeira-Leggett equation is not symmetric in position and momentum the matrix $M$ is not necessarily symmetric. In addition we find that the terms associated with a single power of the operators (left upper $2\times 2$ entries) factorize from the contributions related to products of operators (right bottom $3\times 3$ entries). Due to the fact that the constant terms vanish except for the entry $D_{p^{2}p^{2}}$, we obtain as solution
$c^{(x)}_{T}$ $=0,$ (39)
$c^{(p)}_{T}$ $=0,$
@@
IV Non-equilibrium metrology for a squeezed gaussian state
@@ IV.1 Non-equilibrium Thermometry
$\begin{bmatrix}\begin{array}[]{c}-b\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-2b\langle p^{2}\rangle\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}\\[20.0pt] \begin{array}[]{c}2c\langle x^{2}\rangle\\ +\gamma\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}2\gamma\langle p^{2}\rangle-\dfrac{2\gamma T}{b}\\ +c\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}-4b\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-4b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\end{array}&\begin{array}[]{c}-8b\langle x^{2}\rangle\langle p^{2}\rangle-2b\\ -2b\langle\{x,p\}\rangle^{2}\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}4c\langle x^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\langle\{x,p\}\rangle^{2}\end{array}&\begin{array}[]{c}8\gamma\langle p^{2}\rangle^{2}-\dfrac{8\gamma T}{b}\langle p^{2}\rangle\\ +4c\langle p^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}8c\langle x^{2}\rangle\langle p^{2}\rangle+2c+2c\langle\{x,p\}\rangle^{2}\\ +8\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle-\dfrac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}\\[20.0pt] \begin{array}[]{c}0\end{array}&\begin{array}[]{c}0\end{array}&\begin{array}[]{c}8c\langle x^{2}\rangle^{2}\\ +2b\\ +4\gamma\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}-8b\langle p^{2}\rangle^{2}-2c+4\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2c\langle\{x,p\}\rangle^{2}-\dfrac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}8\gamma\langle x^{2}\rangle\langle p^{2}\rangle-\dfrac{8\gamma T}{b}\langle x^{2}\rangle-2\gamma\\ +8c\langle x^{2}\rangle\langle\{x,p\}\rangle-8b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ +2\gamma\langle\{x,p\}\rangle^{2}\end{array}\end{bmatrix}$ (45)
@@
IV.2 Relaxation rate metrology
@@
Figure 4: Evolution of the Quantum Fisher Information for temperature measurements (green solid) and for relaxation rate measurement (purple solid) for the two cases of $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). Note the qualitatively different behavior with ${\cal F}_{T}$ increasing towards late times, while ${\cal F}_{\gamma}$
@@ V Summary and Conclusion
Acknowledgements.
AKR thanks Korea University for support through project K2503291 Ab-initio simulation of the real-time dynamics of non-relativistic fermions as well as project K2511131. VLP was supported by Xunta de Galicia project ED481A 2022/286, by European Research Council project ERC-2018-ADG-835105 YoctoLHC, by Xunta de Galicia (CIGUS Network of Research Centres), by European Union ERDF, by the Spanish Research State Agency under projects PID2020-119632GBI00 and PID2023-152762NB-I00, and project CEX2023-001318-M financed by MCIN/AEI/10.13039/501100011033
@@ Appendix A Sketch of the derivation of the system matrix M
$\partial_{T}\partial_{t}\hat{\rho}(t,T)$ $=-i[\hat{H},\partial_{T}\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\partial_{T}\hat{\rho}\}]$ (56)
$-2\gamma m[\hat{x},[\hat{x},\hat{\rho}]]-2\gamma mT[\hat{x},[\hat{x},\partial_{T}\hat{\rho}]],$
@@
$\partial_{t}\partial_{T}\hat{\rho}(T)$ $=\frac{1}{2}\left(\hat{\Lambda}_{T}\partial_{t}\hat{\rho}(T)+\partial_{t}\hat{\rho}(T)\hat{\Lambda}_{T}\right)$ (57)
$-\partial_{t}\hat{\rho}(T)\langle\hat{\Lambda}_{T}\rangle-\hat{\rho}(T)\partial_{t}\langle\hat{\Lambda}_{T}\rangle,$
@@
$-\tfrac{1}{2}\left\langle\left\{\hat{x}^{2}\left[\hat{H},\hat{p}^{2}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{x}^{2},\left\{\hat{p},\left[\hat{p}^{2},\hat{x}\right]\right\}\right\}\right\rangle$
$-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{x}^{2},\left[\hat{x},\left[\hat{x},\hat{p}^{2}\right]\right]\right\}\right\rangle+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{x}^{2},\hat{p}^{2}\right\},\hat{x}\right]\right\}\right\rangle$
$+\tfrac{\gamma T}{2b}\left\langle\left[\hat{x},\left[\hat{x},\left\{\hat{x}^{2},\hat{p}^{2}\right\}\right]\right]\right\rangle-i\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x}^{2},\hat{H}\right]\right\rangle$
$-i\gamma\left\langle\hat{p}^{2}\right\rangle\left\langle\left\{\hat{p},\left[\hat{x}^{2},\hat{x}\right]\right\}\right\rangle-\tfrac{\gamma T}{b}\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{x}^{2}\right]\right]\right\rangle.$
@@
$[\hat{x},\hat{p}^{m}]=im\hat{p}^{m-1},\quad[\hat{x}^{n},\hat{p}]=in\hat{x}^{n-1},$
$[\hat{x},\{\hat{x},\hat{p}\}]=2i\hat{x},\quad[\hat{p},\{\hat{x},\hat{p}\}]=-2i\hat{p},$
$[\hat{x}^{2},\hat{p}^{2}]=2i\{\hat{x},\hat{p}\},$

### v2

@@ 1 Introduction
This paper is structured as follows. In section 2 we introduce our formalism to determine the SLD from the master equation. Section 3 applies the formalism to thermometry in the equilibrium steady state of quantum Brownian motion with a harmonic potential, before constructing in section 4 the SLD for both temperature and relaxation rate in a non-equilibrium scenario based on a squeezed Gaussian state.
2 Formalism
@@
3 Equilibrium probe thermometry in Quantum Brownian Motion
@@
$\partial_{t}\hat{\rho}(t,T)=$ (11)
$-i[\hat{H},\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\hat{\rho}\}]-2\gamma mT[\hat{x},[\hat{x},\hat{\rho}]],$
@@
$-\tfrac{1}{2}\left\langle\left\{\hat{A}_{i},\left[\hat{H},\hat{A}_{j}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{A}_{i},\left\{\hat{p},\left[\hat{A}_{j},\hat{x}\right]\right\}\right\}\right\rangle$
$-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{A}_{i},\left[\hat{x},\left[\hat{x},\hat{A}_{j}\right]\right]\right\}\right\rangle$
$+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{A}_{i},\hat{A}_{j}\right\},\hat{x}\right]\right\}\right\rangle$
@@
$-i\gamma\left\langle\hat{A}_{j}\right\rangle\left\langle\left\{\hat{p},\left[\hat{A}_{i},\hat{x}\right]\right\}\right\rangle$
$-\tfrac{\gamma T}{b}\left\langle\hat{A}_{j}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{A}_{i}\right]\right]\right\rangle$
@@
A careful evaluation of all $\times 5$ 5 entries of $M$ yields
$M=$ $\left[\begin{array}[]{ccccc}0&-2b\left\langle p^{2}\right\rangle&0&0&0\\[10.0pt] 2c\left\langle x^{2}\right\rangle&\begin{array}[]{c}2\gamma\left\langle p^{2}\right\rangle\\ -\frac{2\gamma T}{b}\end{array}&0&0&0\\[4.0pt] 0&0&0&2\gamma&-8b\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle-2b\\[4.0pt] 0&0&0&\begin{array}[]{c}8\gamma\left\langle p^{2}\right\rangle^{2}\\ -\frac{8\gamma T\left\langle p^{2}\right\rangle}{b}\end{array}&8c\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle+2c\\[4.0pt] 0&0&\begin{array}[]{c}2b\\ +8c\left\langle x^{2}\right\rangle^{2}\end{array}&\begin{array}[]{c}-8b\left\langle p^{2}\right\rangle^{2}\\ -2c\end{array}&\begin{array}[]{c}-\frac{8\gamma T\left\langle x^{2}\right\rangle}{b}-2\gamma\\ +8\gamma\left\langle p^{2}\right\rangle\left\langle x^{2}\right\rangle\end{array}\\ \end{array}\right]$ (36)
$=$ $\begin{bmatrix}0&-\frac{ba^{2}}{2}&0&0&0\\ \frac{a^{2}b}{2}&-\frac{\gamma a^{2}}{2}&0&0&0\\ 0&0&0&2\gamma&-\dfrac{a^{4}b^{2}}{2c}-2b\\ 0&0&0&-\dfrac{\gamma a^{4}}{2}&\frac{a^{4}b}{2}+2c\\ 0&0&\frac{a^{4}b^{2}}{2c}+2b&-\frac{a^{4}b}{2}-2c&-\frac{\gamma a^{4}b}{2c}-2\gamma\end{bmatrix},$
@@
Interesting structure emerges in the system of linear equations. Since the Caldeira-Leggett equation is not symmetric in position and momentum the matrix $M$ is not necessarily symmetric. In addition we find that the terms associated with a single power of the operators (left upper $\times 2$ 2 entries) factorize from the contributions related to products of operators (right bottom $\times 3$ 3 entries). Due to the fact that the constant terms vanish except for the entry $D_{p^{2}p^{2}}$, we obtain as solution
$c^{(x)}_{T}$ =0=0,= 0 , (38)
$c^{(p)}_{T}$ =0=0,= 0 ,
@@
4 Non-equilibrium metrology for a squeezed gaussian state
@@ 4.1 Non-equilibrium Thermometry
$\begin{bmatrix}\begin{array}[]{c}\scriptstyle-b\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle-2b\langle p^{2}\rangle\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}\\[10.0pt] \begin{array}[]{c}\scriptstyle 2c\langle x^{2}\rangle\\ \scriptstyle+\gamma\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle 2\gamma\langle p^{2}\rangle-\frac{2\gamma T}{b}\\ \scriptstyle+c\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}\\[20.0pt] \begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle-4b\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle-4b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ \scriptstyle+2\gamma\end{array}&\begin{array}[]{c}\scriptstyle-8b\langle x^{2}\rangle\langle p^{2}\rangle-2b\\ \scriptstyle-2b\langle\{x,p\}\rangle^{2}\end{array}\\[20.0pt] \begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 4c\langle x^{2}\rangle\langle\{x,p\}\rangle\\ \scriptstyle+2\gamma\langle\{x,p\}\rangle^{2}\end{array}&\begin{array}[]{c}\scriptstyle 8\gamma\langle p^{2}\rangle^{2}-\frac{8\gamma T}{b}\langle p^{2}\rangle\\ \scriptstyle+4c\langle p^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle 8c\langle x^{2}\rangle\langle p^{2}\rangle+2c+2c\langle\{x,p\}\rangle^{2}\\ \scriptstyle+8\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle-\frac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}\\[20.0pt] \begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 0\end{array}&\begin{array}[]{c}\scriptstyle 8c\langle x^{2}\rangle^{2}\\ +2b\\ \scriptstyle+4\gamma\langle x^{2}\rangle\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle-8b\langle p^{2}\rangle^{2}-2c+4\gamma\langle p^{2}\rangle\langle\{x,p\}\rangle\\ \scriptstyle+2c\langle\{x,p\}\rangle^{2}-\frac{4\gamma T}{b}\langle\{x,p\}\rangle\end{array}&\begin{array}[]{c}\scriptstyle 8\gamma\langle x^{2}\rangle\langle p^{2}\rangle-\frac{8\gamma T}{b}\langle x^{2}\rangle-2\gamma\\ \scriptstyle+8c\langle x^{2}\rangle\langle\{x,p\}\rangle-8b\langle p^{2}\rangle\langle\{x,p\}\rangle\\ \scriptstyle+2\gamma\langle\{x,p\}\rangle^{2}\end{array}\end{bmatrix}$ (44)
@@
4.2 Relaxation rate metrology
@@
Figure 4: Evolution of the Quantum Fisher Information for temperature measurements (green solid) and for relaxation rate measurement (purple solid) for the two cases of $\omega/m=4/3$ (left) and $\omega/m=2/3$ (right). Note the qualitatively different behavior with ${\cal F}_{T}$ increasing towards late times, while ${\cal F}_{\gamma}$ drops to zero.
@@ 5 Summary and Conclusion
We thank Alfonso V. Ramallo for useful discussions. AR thanks Korea University for support through project K2503291 Ab-initio simulation of the real-time dynamics of non-relativistic fermions as well as project K2511131. VLP was supported by Xunta de Galicia project ED481A 2022/286, by European Research Council project ERC-2018-ADG-835105 YoctoLHC, by Xunta de Galicia (CIGUS Network of Research Centres), by European Union ERDF, by the Spanish Research State Agency under projects PID2020-119632GBI00 and PID2023-152762NB-I00, and project CEX2023-001318-M financed by MCIN/AEI/10.13039/501100011033
@@ Appendix A Sketch of the derivation of the system matrix M
$\scriptstyle\hat{x}$ $\scriptstyle\hat{p}$ $\scriptstyle\hat{x}^{2}$ $\scriptstyle\hat{x}\hat{p}$ $\scriptstyle\hat{p}^{2}$ $\scriptstyle\hat{p}\hat{x}$
$\scriptstyle\hat{x}^{2}$ $\scriptstyle\hat{x}^{3}$ $\scriptstyle W(\hat{x}^{2}\hat{p})+i\hat{x}$ $\scriptstyle\hat{x}^{4}$ $\scriptstyle W(\hat{x}^{3}\hat{p})+\frac{3i}{2}\hat{x}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})+2iW(\hat{x}\hat{p})-\frac{1}{2}$ $\scriptstyle W(\hat{x}^{3}\hat{p})+\frac{i}{2}\hat{x}$
$\scriptstyle\hat{x}\hat{p}$ $\scriptstyle W(\hat{x}^{2}\hat{p})$ $\scriptstyle W(\hat{x}\hat{p}^{2})+i\hat{p}$ $\scriptstyle W(\hat{x}^{3}\hat{p})-\frac{i}{2}\hat{x}^{2}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})+iW(\hat{x}\hat{p})$ $\scriptstyle W(\hat{x}\hat{p}^{3})+\frac{3i}{2}\hat{p}^{2}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})+\frac{1}{2}$
$\scriptstyle\hat{p}^{2}$ $\scriptstyle W(\hat{x}\hat{p}^{2})-i\hat{p}$ $\scriptstyle\hat{p}^{3}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})-2iW(\hat{x}\hat{p})-\frac{1}{2}$ $\scriptstyle W(\hat{x}\hat{p}^{3})-\frac{i}{2}\hat{p}^{2}$ $\scriptstyle\hat{p}^{4}$ $\scriptstyle W(\hat{x}\hat{p}^{3})-\frac{3i}{2}\hat{p}^{2}$
$\scriptstyle\hat{p}\hat{x}$ $\scriptstyle W(\hat{x}^{2}\hat{p})-i\hat{x}$ $\scriptstyle W(\hat{x}\hat{p}^{2})$ $\scriptstyle W(\hat{x}^{3}\hat{p})-\frac{3i}{2}\hat{x}^{2}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})+\frac{1}{2}$ $\scriptstyle W(\hat{x}\hat{p}^{3})+\frac{i}{2}\hat{p}^{2}$ $\scriptstyle W(\hat{x}^{2}\hat{p}^{2})-iW(\hat{x}\hat{p})$
Table 1: Weyl correspondences for every combination of three or four operators appearing in the derivation of the matrix elements $M_{ij}$. The operator on the left column times the operator on the top row is equal to the corresponding cell, e.g.: $\hat{x}\hat{p}^{2}\hat{x}=(\hat{x}\hat{p})(\hat{p}\hat{x})=W(\hat{x}^{2}\hat{p}^{2})+\frac{1}{2}$
@@
$\partial_{T}\partial_{t}\hat{\rho}(t,T)$ $=-i[\hat{H},\partial_{T}\hat{\rho}]-i\gamma[\hat{x},\{\hat{p},\partial_{T}\hat{\rho}\}]-2\gamma m[\hat{x},[\hat{x},\hat{\rho}]]-2\gamma mT[\hat{x},[\hat{x},\partial_{T}\hat{\rho}]],$ (55)
@@
$\partial_{t}\partial_{T}\hat{\rho}(T)$ $=\frac{1}{2}\left(\hat{\Lambda}_{T}\partial_{t}\hat{\rho}(T)+\partial_{t}\hat{\rho}(T)\hat{\Lambda}_{T}\right)-\partial_{t}\hat{\rho}(T)\langle\hat{\Lambda}_{T}\rangle-\hat{\rho}(T)\partial_{t}\langle\hat{\Lambda}_{T}\rangle,$ (56)
@@
$-\tfrac{1}{2}\left\langle\left\{\hat{x}^{2},\left[\hat{H},\hat{p}^{2}\right]\right\}\right\rangle-\tfrac{i}{2}\left\langle\left\{\hat{x}^{2},\left\{\hat{p},\left[\hat{p}^{2},\hat{x}\right]\right\}\right\}\right\rangle-\tfrac{\gamma T}{2b}\left\langle\left\{\hat{x}^{2},\left[\hat{x},\left[\hat{x},\hat{p}^{2}\right]\right]\right\}\right\rangle+\tfrac{i\gamma}{2}\left\langle\left\{\hat{p},\left[\left\{\hat{x}^{2},\hat{p}^{2}\right\},\hat{x}\right]\right\}\right\rangle$
$+\tfrac{\gamma T}{2b}\left\langle\left[\hat{x},\left[\hat{x},\left\{\hat{x}^{2},\hat{p}^{2}\right\}\right]\right]\right\rangle-i\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x}^{2},\hat{H}\right]\right\rangle-i\gamma\left\langle\hat{p}^{2}\right\rangle\left\langle\left\{\hat{p},\left[\hat{x}^{2},\hat{x}\right]\right\}\right\rangle-\tfrac{\gamma T}{b}\left\langle\hat{p}^{2}\right\rangle\left\langle\left[\hat{x},\left[\hat{x},\hat{x}^{2}\right]\right]\right\rangle.$
@@
$[\hat{x},\hat{p}^{m}]=im\hat{p}^{m-1},\quad[\hat{x}^{n},\hat{p}]=in\hat{x}^{n-1},\quad[\hat{x},\{\hat{x},\hat{p}\}]=2i\hat{x},\quad[\hat{p},\{\hat{x},\hat{p}\}]=-2i\hat{p},\quad[\hat{x}^{2},\hat{p}^{2}]=2i\{\hat{x},\hat{p}\},$